HIGH-RESOLUTION CAVITY ENHANCED ABSORPTION SPECTROSCOPY OF CARBON DIOXIDE AND METHANE AT 1.5 μm

TONGMEI MA, HONGBING CHEN, \underline{A. S-C. CHEUNG}, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.